Md lab needs to make 500 gallons of 34% acid solution. The only solutions avalible are 25% solution and a 50% solution. Tow many gallons of each solution should be mixed to make the 34% solution
Answer by mananth(16946)

percent ---------------- quantity
Acid type I 25.00% ---------------- x gallons
Acid type II 50.00% ------ 500 - x gallons
Mixture 34.00% ---------------- 500
500
25.00% x + 50.00% ( 500 - x ) = 34.00% * 500
25 x + 50 ( 500 - x ) = 17000
25 x + 25000 - 50 x = 17000
25 x - 50 x = 17000 - -25000
-25 x = -8000
/ -25
x = 320 gallons 25.00% Acid type I
180 gallons 50.00% Acid type II

m.ananth@hotmail.ca